During the annual Black Friday Sale, The OLX sold a pair of ski boots, regularly priced at $245.00, at a discount of 40%. The boots cost $96.00 and expenses are 26% of the regular selling price. For how much were the ski boots sold?

1
Expert's answer
2021-03-10T06:32:54-0500

1)Here we compute the discount first which is:

(40/100)*245=98.00

2)After that we compute the (Selling Price -Discount) to get the Buying Price hence,

245-98=147

total expenses include 26% of regular price.

(26/100)*245=63.70

Therefore total Selling price is :

147+63.70=210.70

$210